OpenStudy (anonymous):

write the absolute value function f(x)=|x+4| as a piecewise function.

OpenStudy (anonymous):

f(x)= -(x+4) , x<-4 =(x+4) , x>-4

OpenStudy (anonymous):

f(x)= -(x+4) , x<-4 =(x+4) , x>= -4